# GPU-Accelerated Algorithms for Process Mapping thanks: Henning Woydt is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – DFG SCHU 2567/6-1

### v1

@@ Abstract
Process mapping asks to assign vertices of a task graph to processing elements of a supercomputer such that the computational workload is balanced while the communication cost is minimized. Motivated by the recent success of GPU-based graph partitioners, we propose two GPU-accelerated algorithms for this optimization problem.
The first algorithm employs hierarchical multisection, which partitions the task graph alongside the hierarchy of the supercomputer. The method utilizes GPU-based graph partitioners to accelerate the mapping process. The second algorithm integrates process mapping directly into the modern multilevel graph partitioning pipeline. Vital phases like coarsening and refinement are accelerated by exploiting the parallelism of GPUs.
In our experiments, both methods achieve speedups exceeding 300 when compared to state-of-the-art CPU-based algorithms. The first algorithm has, on average, about 10 percent greater communication costs and thus remains competitive to CPU algorithms. The second approach is much faster, with a geometric mean speedup of 77.6 and peak speedup of 598 at the cost of lower solution quality. To our knowledge, these are the first GPU-based algorithms for process mapping.
I Introduction
@@
Parallel high-performance applications rely on the efficient execution of large numbers of tasks across massively parallel architectures. This involves placing individual tasks on the Processing Elements (PEs) of the underlying hardware architecture such that the computational workload is evenly balanced. However, load balancing alone is not sufficient. In parallel applications, the individual tasks usually communicate with each other, i.e., they send and receive data to and from each other. The cost of this communication heavily depends on the location of the tasks in hardware topology. Modern supercomputers are typically organized hierarchically, for example into processors, nodes, racks and islands. The communication cost between two PEs, therefore, depends on their relative locations within this hierarchy. Two PEs on the same processor communicate data much more quickly than two PEs which are located on different islands. Thus, in addition to balancing the workload, it is advantageous to minimize the total communication cost by placing highly interactive tasks close to each other. This optimization problem is known as the general process mapping problem (GPMP).
@@
Optimally solving the GPMP is $\mathrm{NP}$-hard [39] as it involves solving the well-known quadratic assignment problem (QAP) [9]. The QAP asks for a set of $n$ facilities and $n$ locations to minimize the total cost of assigning each facility to one location, where the cost depends on both the flow between facilities and the distance between their locations. Since the QAP is strongly $\mathrm{NP}$-hard, i.e., no constant factor approximation exists unless $\mathrm{P}=\mathrm{NP}$, its hardness transfers to the GPMP. The hardness is also reflected in practice as no meaningful instance with $n>20$ can be solved in a reasonable time [9]. Moreover, the same hardness result arises from graph partitioning (GP), where the goal is to partition a graph while minimizing the number of edges between them. Graph partitioning is $\mathrm{NP}$-complete and also admits no constant-factor approximation unless $\mathrm{P}=\mathrm{NP}$.
@@
The GPMP is commonly solved using either the two-phase approach or via integrated mapping. In the two-phase approach, the task set is first partitioned into as many (almost) equally sized sets as there are PEs in the supercomputer. While partitioning the task set, the necessary communication between the resulting sets is minimized. In the second phase, the subsets are mapped to the PEs such that the total cost of communication is minimized, taking the distances between the PEs into account. The first phase is an instance of GP, while second phase is a QAP instance. The two-phase approach is often easier to realize as it can fall back to two well-studied problems, i.e., graph partitioning the QAP. The algorithm SharedMap [44] currently offers state-of-the-art solution quality by utilizing hierarchical multisection, a special variant of the two-phase approach.
The integrated mapping approach combines both phases into a single process by incorporating the mapping objective directly into the graph partitioning pipeline. Rather than minimizing the number of edges crossing between the subsets, as in classical graph partitioning, it directly minimizes the total cost of communication. Typically, this is achieved using the modern multilevel framework, which involves coarsening the input graph to a smaller representation, computing an initial mapping, and then refining it during uncoarsening. The approach is more complex to realize as the whole graph partitioning pipeline has to be adopted to the new objective function. However, this also enables function specific optimization, which result in an overall faster execution time. As a result, integrated mapping often achieves faster runtimes compared to the two-phase approach.
@@
Recently, Graphic Processing Units (GPUs) have been utilized to tackle graph partitioning. GPUs provide massive parallelism that can significantly reduce runtime, when exploited effectively. However, leveraging thousands of GPU cores is challenging due to the irregularity of graphs and lack of straightforward parallelization strategies. Despite this, recent GPU-based algorithms like Jet [18] and G-kway [31] have demonstrated promising results. They deliver solution qualities comparable to the serial CPU-based partitioner Metis [27], while being up to 20 times faster. Since both the two-phase and integrated mapping approach rely on graph partitioning, these results should transfer to GPMP.
[…]
II Preliminaries
II-A Concepts and Notations
@@
Let $G=(V,E)$ with $n=|V|$ and $m=|E|$ be an undirected graph, and let $c:V\mapsto\mathbb{R}$ and $\omega:E\mapsto\mathbb{R}$ be vertex and edge weights. Their extension to sets is $c(V^{\prime})=\sum_{v\in V^{\prime}}c(v)$ and $\omega(E^{\prime})=\sum_{e\in E^{\prime}}\omega(e)$ for vertex sets $V^{\prime}$ and edge sets $E^{\prime}$.
@@
The graph partitioning problem (GPP) asks for a partition of the vertex set $V$ into $k$ equally sized sets. More formally let $k\in\mathbb{N}$ be the desired number of blocks and let $\epsilon\geq 0$ be a hyper-parameter called imbalance. GPP then asks for a partition $V=V_{1}\cup\ldots\cup V_{k}$ such that $\forall i\neq j:V_{i}\cap V_{j}=\emptyset$ and each $V_{i}$ adheres to the balance constraint, that is $c(V_{i})\leq L_{\max}\coloneqq(1+\epsilon)\frac{c(V)}{k}$. Such a partition is called a $k$-way partition of $G$. Let $E_{ij}=\{\{u,v\}\in E\mid u\in V_{i},v\in V_{j}\}$ be the edges crossing between block $V_{i}$ and $V_{j}$. The edge-cut of a partition is the total weight of edges crossing between blocks, i.e., $\sum_{i<j}\omega(E_{ij})$. The goal of GP is to determine a balanced $k$-way partition that minimizes the edge-cut.
@@
A parallel application consists of $n$ weighted tasks which communicate with each other. This is captured by the communication matrix $\mathcal{C}\in\mathbb{R}^{n\times n}$, where an entry $\mathcal{C}_{ij}$ denotes the communication volume between tasks $i$ and $j$. An entry $\mathcal{C}_{ij}=0$ denotes that the tasks do not communicate with each other. Most modern parallel applications only have a very sparse communication pattern between the tasks, which is why we model the communication matrix $\mathcal{C}$ as a task graph $G_{\mathcal{C}}$. For each non-zero entry $\mathcal{C}_{ij}$ the task graph has one forward and backward edge with weight $\mathcal{C}_{ij}$.
@@
Let $k$ be the number of PEs on a supercomputer. The cost of communication between the PEs is captured by the hardware topology matrix $\mathcal{D}\in\mathbb{R}^{k\times k}$. An entry $\mathcal{D}_{xy}$ denotes a multiplicative cost factor that is applied to communication between PEs $x$ and $y$. The cost of task $i$ and $j$ communicating, if placed on PEs $x$ and $y$ respectively, is $\mathcal{C}_{ij}\mathcal{D}_{xy}$.
@@
The focus of this work is to tackle the general process mapping problem. It asks to assign the vertices of the task graph $G_{\mathcal{C}}$ to PEs of the supercomputer such that the amount of work on each PE is balanced and the total cost of communication is minimized. More formally, we ask for a mapping $\Pi:[n]\mapsto[k]$ such that $J(C,D,\Pi)\coloneqq\sum_{i,j}\mathcal{C}_{ij}\mathcal{D}_{\Pi(i)\Pi(j)}$ is minimized, while adhering to the balance constraint.
@@
In the hierarchical process mapping problem (HPMP), a special variant of the GPMP, the hardware topology of the supercomputer is described by a homogeneous hierarchy $H=a_{1}:\ldots:a_{\ell}$. It denotes that each processor has $a_{1}$ PEs, each node has $a_{2}$ processors, each rack has $a_{3}$ nodes, and so forth. In total, the system has $k=\prod_{1\leq\ell}a_{i}$ PEs. Corresponding to the hierarchy, a distance $D=d_{1}:\ldots:d_{\ell}$ is given which denotes the cost factor between the PEs. Two PEs on the same processor have a cost factor of $d_{1}$, two PEs on the same node but on different processors have a cost factor of $d_{2}$ and so forth.
@@
In GPMP we typically assume $n>k$, i.e., we have (considerably) more tasks than PEs. If $n=k$ the problem is known as the one-to-one process mapping problem (OPMP) which is identical to the quadratic assignment problem (QAP). The QAP is commonly solved using refactorization and linearization techniques [3, 2, 21], of its corresponding mixed integer linear programming formulation [30]. We refer the reader to [9, 4, 32] for comprehensive surveys. The problems GPP, GPMP, HPMP and QAP are all NP-hard problems [17, 39] and no constant factor approximation guarantee exists unless P = NP.
II-B Multilevel Graph Partitioning
@@
Modern graph partitioning relies on the multilevel approach, which works in three phases, coarsening, partitioning and uncoarsening with refinement. In the coarsening phase a succession of smaller graphs $G,G_{1},\ldots,G_{t}$ is computed, with the goal that a smaller graph $G_{i+1}$ retains the same overall structure as $G_{i}$. A “good” partition of $G_{i+1}$ would then also resemble a “good” partition for $G_{i}$. Coarsening is applied until the coarsest graph has less than $ck$ vertices, where $c$ is a tunable constant. To coarsen a graph, one usually applies a matching or a clustering algorithm. The matching algorithm rates each edge and then determines a maximum weighted matching. Edge contraction is then applied on the determined matching. Edge contraction on an edge $\{u,v\}$ is applied by replacing vertices $u$ and $v$ with a new vertex $w$ with weight $c(w)=c(u)+c(v)$. Former neighbors of $u$ and $v$ are connected to $w$ with the same edge weight. Parallel edges are fused and their weight is summed. A clustering algorithm first determines clusters in the graph, and then contracts each cluster into one super-vertex. The vertex weights and parallel edges are handled in the same way as in the matching.
@@
In the initial partitioning phase, the coarsest graph $G_{t}$ is partitioned. Since $G_{t}$ is very small, expensive but high-quality algorithms, such as recursive bisection or direct $k$-way partitioning, can be applied to minimize edge-cut.
@@
The uncoarsening and refinement phase iteratively uncontracts the graphs $G_{i+1}$ into $G_{i}$ by undoing the contractions. After each uncontraction, local search algorithms are applied to improve the objective. The algorithms try to move vertices across the blocks to minimize the edge-cut, while respecting the balance constraint. Widely used refinement algorithms are the Fiduccia-Mattheyses (FM) algorithm [16], label propagation [38] and flow-based refinement [40]. Notable serial graph partitioners are Metis [27], KaFFPa [40], Jostle [48], Scotch [36] and KaHyPar [41].
III Related Work
@@
Process mapping is closely related to graph partitioning, which has already seen a tremendous amount of research over the last few decades. We already gave a brief overview in Section II-B and refer the reader to [6, 10, 8, 42] for more comprehensive surveys. We will here give an overview of Jet [18], a GPU-based graph partitioning (Section III-A), and of process mapping (Section III-B).
III-A GPU Multilevel Graph Partitioning
@@
Gilbert et al. [18] present Jet, a GPU-accelerated multilevel graph partitioner designed to achieve high partition quality and fast runtimes. Their main contribution includes a novel GPU-parallel refinement algorithm based on label propagation [38]. They implemented their algorithm in the performance-portable Kokkos framework [14], enabling execution on GPUs and multi-core CPUs with a single code base.
@@
Matching. Jet utilizes a two-hop matching [29], originally developed for Mt-Metis [28], a shared-memory graph partitioner. The algorithm first applies a standard heavy-edge matching and only if more than 25% of vertices remain unmatched a two hop-matching is applied. A heavy edge-matching determines for each vertex its heaviest edge and two neighbors are matched if they agree on their heaviest edge. The two-hop matching can be split into three categories: leaf, twin and relative vertices. Leaf vertices have degree one, two twin vertices share the same neighborhood and relative vertices share at least one common neighbor. To identify twin vertices the neighborhoods of all vertices are hashed and sorted, so vertices with the same neighborhood can easily be found. To accelerate the matching of relatives, special matchmaker vertices with small degree are used. The two-hop matching first matches leaf vertices that share a neighbor, then twin vertices and in the end relative vertices, if necessary.
@@
Coarsening. To coarsen the graph, a fine-grained per vertex hashing algorithm is deployed. First the maximum neighborhood size of each vertex in the coarse graph is overestimated. Then a prefix-sum over the sizes is determined, assigning each vertex a subarray to store its neighbors. Each vertex hashes its new neighboring vertices and inserts them and their weight into their respective subarray. In the end, the edges of all vertices are extracted into the classic Compressed Sparse Row (CSR) format.
@@
Gilbert et. al [18] tackle these problems by splitting the refinement into two parts, first unconstrained label propagation and afterwards rebalancing. The unconstrained label propagation starts by determining for each vertex $v$ the gain $F(v)$ (decrease in edge-cut) of moving it in an adjacent block. All potential vertex moves are then passed through the first filter $F(v)\geq 0\lor-F(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. Here, $\operatorname{conn}(v,b)$ is the sum of edge weights from vertex $v$ to block $b$. To pass the first filter a move needs to be non-negative (fist condition) or its increase in edge-cut must be comparatively small (second condition). The constant $c\in[0,1]$ controls the number of negative gain moves. These moves are allowed to escape local minima. All moves that pass the first filter are collected in a list $\mathcal{X}$, which is implicitly sorted decreasingly by gain. For each vertex $v\in\mathcal{X}$, the algorithm then recomputes the gain $F_{2}(v)$ under an approximate future partition state: if a neighbor $u$ appears earlier in the ordering, it is assumed to have already moved to its new block. In a second filter only the vertices $v\in\mathcal{X}$ with a non-negative gain $F_{2}(v)$ are moved to their respective destination. To prevent oscillation each vertex that was moved is locked in the next iteration of unconstrained label propagation.
@@
Rebalancing. Since the unconstrained label propagation phase ignores the balance constraint, an additional rebalancing phase is required. Jet implements two similar strategies, weak and strong rebalancing. Weak rebalancing determines for each vertex in an oversized partition $a$ the loss of moving it to a valid destination block. Valid destination blocks are blocks $b$ with $c(b)\leq\sigma\leq L_{\max}$. The upper bound $\sigma$ is used instead of $L_{\max}$ to create a small buffer zone, so the block is not overloaded after only one move. For each oversized block $a$ its vertices are approximately sorted increasingly by loss in a list $L_{a}$. Then a prefix $L_{a}^{*}\subseteq L_{a}$ is determined such that moving all vertices in $L_{a}^{*}$ will balance block $a$. This minimizes the loss to balance a specific block, however, the destination blocks might become oversized, so another iteration might be necessary. Strong rebalancing reverses the process. Instead of oversized blocks pushing their vertices out, underloaded blocks will pull vertices from oversized blocks. They will only pull so many vertices such they themselves will not become oversized. This often achieves a balanced partition in just one iteration, however the loss is typically greater as in weak rebalancing. In practice, Jet applies two iterations of weak rebalancing followed by one iteration of strong rebalancing if necessary.
@@
In total, Jet combines its unconstrained label propagation and rebalancing in multiple iterations and keeps the best solution encountered. The algorithm offers comparable solution quality to the well established CPU-based graph partitioner Metis [27], while being up to 20 times faster. Other GPU graph partitioners are Goodarzi et al. [19], Sphynx [1] and G-kway [31].
III-B Process Mapping

### v2

@@ Abstract
Process mapping asks to assign vertices of a task graph to processing elements of a supercomputer such that the computational workload is balanced while the communication cost is minimized. Motivated by the recent success of GPU-based graph partitioners, we propose two GPU-accelerated algorithms for this optimization problem. The first algorithm employs hierarchical multisection, which partitions the task graph alongside the hierarchy of the supercomputer. The method utilizes GPU-based graph partitioners to accelerate the mapping process. The second algorithm integrates process mapping directly into the modern multilevel graph partitioning pipeline. Vital phases like coarsening and refinement are accelerated by exploiting the parallelism of GPUs. The first algorithm has, on average, about 12 percent higher communication costs than the state-of-the-art solver and thus remains competitive with it. However, in terms of speed, it vastly outperforms the competitor with a geometric mean speedup of 22 times and a maximum speedup of 934 times. The second approach is even faster, with a geometric mean speedup of 1454 times and a peak speedup of 12376 times. Compared to other algorithms that prioritize speed over solution quality, this approach has the same quality but much greater speedups. To our knowledge, these are the first GPU-based algorithms for process mapping.
keywords:
GPU, Process Mapping, Graph Partitioning
1 Introduction
@@
Parallel high-performance applications rely on the efficient execution of large numbers of tasks across massively parallel architectures. This involves placing individual tasks on the Processing Elements (PEs) of the underlying hardware architecture such that the computational workload is evenly balanced. However, load balancing alone is not sufficient. In parallel applications, the individual tasks usually communicate with each other, i.e., they send and receive data to and from each other. The cost of this communication heavily depends on the location of the tasks in the hardware topology. Modern supercomputers are typically organized hierarchically, for example into processors, nodes, racks, and islands. The communication cost between two PEs, therefore, depends on their relative locations within this hierarchy. Two PEs on the same processor communicate data much more quickly than two PEs that are located on different islands. Thus, in addition to balancing the workload, it is advantageous to minimize the total communication cost by placing highly interactive tasks close to each other. This optimization problem is known as the general process mapping problem (GPMP).
@@
Optimally solving the GPMP is $\mathrm{NP}$-hard [40] as it involves solving the well-known quadratic assignment problem (QAP) [10]. The QAP considers a set of $n$ facilities and $n$ locations. The objective is to assign each facility to exactly one location in a way that minimizes the total cost. This cost is determined by the flow between pairs of facilities and the distance between the locations to which they are assigned. In other words, facilities with higher interaction should ideally be placed closer together to reduce the overall cost. Since the QAP is strongly $\mathrm{NP}$-hard, i.e., no constant factor approximation exists unless $\mathrm{P}=\mathrm{NP}$, its hardness transfers to the GPMP. The hardness is also reflected in practice as no meaningful instance with $n>20$ can be solved in a reasonable time [10]. Moreover, the same hardness result arises from graph partitioning (GP). GP considers a graph $G=(V,E)$ and asks for a partition of the vertex set into $k$ pairwise disjoint sets of equal size, such that the number of edges crossing between the different subsets is minimized. Graph partitioning is $\mathrm{NP}$-complete [18] and also admits no constant-factor approximation unless $\mathrm{P}=\mathrm{NP}$ [4].
@@
The GPMP is commonly solved using either the two-phase approach or via integrated mapping. In the two-phase approach, the task set is first partitioned into as many (almost) equally sized sets as there are PEs in the supercomputer. While partitioning the task set, the necessary volume of communication across the different resulting sets is minimized. In the second phase, the subsets are mapped to the PEs such that the total cost of communication is minimized, taking the distances between the PEs into account. While the first phase is an instance of graph partitioning, the second phase is an instance of the quadratic assignment problem. This makes the approach easier to realize as one can fall back to two well-studied problems. The algorithm SharedMap [45] currently offers state-of-the-art solution quality by utilizing hierarchical multisection. This special variant of the two-phase approach directly exploits the hardware topology of the supercomputer. The integrated mapping approach combines both phases into a single process. The mapping objective of the second phase is directly integrated into the partitioning of the first phase. Rather than minimizing the communication volume between the subsets, it directly minimizes the total cost of communication. Typically, this is achieved using the modern multilevel graph partitioning framework. The input graph is first coarsened to a smaller representation and then an initial partition is determined. During uncoarsening, the partition is refined using local search algorithms. The approach is more complex as the whole graph partitioning pipeline has to be adopted to the new objective function. However, this also enables function-specific optimization resulting in an overall faster running time. Integrated mapping algorithms, like IntMap [16], often achieve faster running times compared to the two-phase approach.
@@
Recently, Graphics Processing Units (GPUs) have been utilized to tackle graph partitioning. GPUs provide many more cores than CPUs and, when exploited efficiently, can significantly reduce the running time. However, effectively leveraging this level of parallelism for graph algorithms is challenging. Graph computations typically involve irregular memory access patterns, uneven workloads across vertices, and complex data dependencies. Despite this, recent GPU-based algorithms like Jet [19] and G-kway [32] have demonstrated promising results. They deliver solution qualities comparable to the serial CPU-based graph partitioner Metis [28] while being up to 20 times faster. Both the two-phase and integrated mapping approach rely on graph partitioning, so these results should transfer to GPMP.
Our Contribution. We present two GPU-based algorithms to solve the process mapping problem. The first one is based on the two-phase approach and utilizes hierarchical multisection. In contrast, the second algorithm integrates the mapping into the graph partitioning pipeline. Both significantly outperform state-of-the-art CPU solvers in runtime, achieving speedups of more than $900\times$, while remaining competitive with their solution quality.
2 Concepts and Notations
@@
Let $G=(V,E)$ with $n=|V|$ and $m=|E|$ be an undirected graph, and let $c:V\mapsto\mathbb{R}$ and $\omega:E\mapsto\mathbb{R}$ be vertex and edge weights. Their natural extension to sets is $c(V^{\prime})=\sum_{v\in V^{\prime}}c(v)$ and $\omega(E^{\prime})=\sum_{e\in E^{\prime}}\omega(e)$ for vertex sets $V^{\prime}\subseteq V$ and edge sets $E^{\prime}\subseteq E$.
@@
The graph partitioning problem (GPP) asks for a partition of the vertex set $V$ into $k$ equally sized sets. More formally, let $k\in\mathbb{N}$ be the desired number of blocks and let $\epsilon\geq 0$ be a hyper-parameter called imbalance. GPP then asks for a partition $V=V_{1}\cup\ldots\cup V_{k}$ such that $\forall i\neq j:V_{i}\cap V_{j}=\emptyset$ and each $V_{i}$ adheres to the balance constraint, that is, $c(V_{i})\leq L_{\max}\coloneqq\lceil(1+\epsilon)\frac{c(V)}{k}\rceil$. Such a partition is called a $k$-way partition of $G$. We call a block $V_{i}$ overloaded if $c(V_{i})>L_{\max}$ and underloaded if $c(V_{i})\leq L_{\max}$. Let $E_{ij}=\{\{u,v\}\in E\mid u\in V_{i},v\in V_{j}\}$ be the edges crossing between block $V_{i}$ and $V_{j}$. The edge-cut of a partition is the total weight of edges crossing between blocks, i.e., $\sum_{i<j}\omega(E_{ij})$. The goal of GP is to determine a balanced $k$-way partition that minimizes the edge-cut.
@@
A parallel application consists of $n$ weighted tasks which communicate with each other. This is captured by the communication matrix $\mathcal{C}\in\mathbb{R}^{n\times n}$, where an entry $\mathcal{C}_{ij}$ denotes the communication volume between tasks $i$ and $j$. An entry $\mathcal{C}_{ij}=0$ denotes that the tasks do not communicate with each other. Most modern parallel applications only have a very sparse communication pattern, which is why the application can also be modeled as a task graph $G_{\mathcal{C}}$. For each task in the application there exists one vertex with the same weight. For each non-zero entry $\mathcal{C}_{ij}$, the task graph has one forward and backward edge with weight $\mathcal{C}_{ij}$. It is assumed that the communication is symmetric, otherwise it can be modeled symmetrically [8].
@@
Let $k$ be the number of PEs on a supercomputer. The cost of communication between two PEs is captured by the hardware topology matrix $\mathcal{D}\in\mathbb{R}^{k\times k}$. An entry $\mathcal{D}_{xy}$ denotes a cost factor that is applied to communication between PEs $x$ and $y$. The cost of tasks $i$ and $j$ communicating, if placed on PEs $x$ and $y$ respectively, is therefore $\mathcal{C}_{ij}\mathcal{D}_{xy}$. It is assumed that all PEs can communicate with each other and that the matrix is symmetric.
@@
The general process mapping problem (GPMP) is then defined as follows. Let $G_{\mathcal{C}}=(V,E)$ with $n=|V|$ be a task graph, let $\mathcal{D}$ be the hardware topology matrix of a supercomputer with $k$ PEs and let $\epsilon$ be the allowed imbalance. GPMP then asks to partition the task set $V$ into $k$ equally sized subsets, such that the overall communication cost is minimized. More formally, determine a mapping $\Pi:[n]\mapsto[k]$ such that $J(\mathcal{C},\mathcal{D},\Pi)\coloneqq\sum_{i,j}\mathcal{C}_{ij}\mathcal{D}_{\Pi(i)\Pi(j)}$ is minimized, while each block adheres to the balance constraint.
@@
In the hierarchical process mapping problem (HPMP), a special variant of the GPMP, the hardware topology of the supercomputer is described by a hierarchy $H=a_{1}:\ldots:a_{\ell}$. It denotes that each processor has $a_{1}$ PEs, each node has $a_{2}$ processors, each rack has $a_{3}$ nodes, and so forth. In total, the system has $k=\prod_{i\leq\ell}a_{i}$ PEs. Corresponding to the hierarchy, a distance $D=d_{1}:\ldots:d_{\ell}$ is given which denotes the cost factor between the PEs. Two PEs on the same processor have a cost factor of $d_{1}$, two PEs on the same node but on different processors have a cost factor of $d_{2}$, two PEs on the same rack but on different islands have a cost factor of $d_{3}$, and so forth. The focus of this work lies in solving the HPMP.
@@
In GPMP, $n>k$ is typically assumed, i.e., there are (considerably) more tasks than PEs. If $n=k$, the problem is known as the one-to-one process mapping problem (OPMP) which is identical to the quadratic assignment problem (QAP). The QAP is commonly solved using refactorization and linearization techniques [3, 2, 22] of its corresponding mixed integer linear programming formulation [31]. We refer the reader to [10, 5, 33] for comprehensive surveys. The problems GPP, GPMP, HPMP, and QAP are all NP-hard problems [18, 40] and no constant factor approximation guarantee exists unless P = NP.
2.1 Multilevel Graph Partitioning
@@
Modern graph partitioning relies on the multilevel approach, which works in three phases: coarsening, partitioning, and uncoarsening with refinement. In the coarsening phase, a succession of smaller graphs $G,G_{1},\ldots,G_{t}$ is computed, with the goal that a smaller graph $G_{i+1}$ retains the same overall structure as $G_{i}$. A “good” partition of $G_{i+1}$ would then also resemble a “good” partition for $G_{i}$. Coarsening is applied until the coarsest graph has fewer than $ck$ vertices, where $c$ is a tunable constant. This small graph can then be efficiently partitioned. The contractions are then undone and the quality of the partition is improved.
To coarsen a graph, one usually applies a matching or a clustering algorithm. The matching algorithm rates each edge and then determines a maximum weighted matching. The edges selected in the matching are then contracted. Contracting an edge $\{u,v\}$ replaces the vertices $u$ and $v$ with a new vertex $w$ whose weight is defined as $c(w)=c(u)+c(v)$. All former neighbors of $u$ and $v$ are subsequently connected to $w$. Parallel edges are fused and their weights are summed. A clustering algorithm first determines clusters in the graph, and then contracts each cluster into one super-vertex. The vertex weights and parallel edges are handled in the same way as in the matching.
@@
In the initial partitioning phase, the coarsest graph $G_{t}$ is partitioned. Since $G_{t}$ is very small, expensive but high-quality algorithms, such as recursive bisection or direct $k$-way partitioning, can be applied to minimize the edge-cut.
@@
The uncoarsening and refinement phase iteratively uncontracts the graphs $G_{i+1}$ into $G_{i}$ by undoing the contractions. After each uncontraction, local search algorithms are applied to improve the objective. The algorithms try to move vertices across the blocks to minimize the edge-cut, while respecting the balance constraint. Widely used refinement algorithms are the Fiduccia-Mattheyses (FM) algorithm [17], label propagation [39], and flow-based refinement [41]. Notable serial graph partitioners are Metis [28], KaFFPa [41], Jostle [49], Scotch [37] and KaHyPar [42].
3 Related Work
@@
Process mapping is closely related to graph partitioning, which has already seen a tremendous amount of research over the last few decades. We already gave a brief overview in Section 2.1 and refer the reader to [7, 11, 9, 43] for more comprehensive surveys. We will give here an overview of Jet [19], a GPU-based graph partitioner, and of process mapping.
3.1 GPU Multilevel Graph Partitioning
@@
Gilbert et al. [19] present Jet, a GPU-accelerated multilevel graph partitioner designed to achieve high partition quality and fast runtimes. Their main contribution includes a novel GPU-parallel refinement algorithm based on label propagation [39]. They implemented their algorithm in the performance-portable Kokkos framework [15].
@@
Matching. Jet utilizes a two-hop matching [30], originally developed for Mt-Metis [29], a shared-memory graph partitioner. The algorithm first applies a heavy-edge matching and only if more than 25% of vertices remain unmatched, a two-hop matching is applied. A heavy-edge matching determines for each vertex its heaviest edge and two neighbors are matched if they agree on their heaviest edge. The two-hop matching can be split into three categories: leaf, twin and relative vertices. Leaf vertices have degree one, two twin vertices share the same neighborhood and relative vertices share at least one common neighbor. To identify twin vertices the neighborhoods of all vertices are hashed and sorted, so vertices with the same neighborhood can easily be found. To accelerate the matching of relatives, special matchmaker vertices with small degree are used. At first leaf vertices that share a neighbor are matched, then twin vertices and at the end, relative vertices, if necessary.
@@
Coarsening. To coarsen the graph, a fine-grained per-vertex hashing algorithm is deployed. First, the maximum neighborhood size of each coarse vertex is overestimated. Then a prefix-sum over the sizes is determined, assigning each vertex a subarray. Each vertex hashes its new neighboring vertices and inserts them and their weights into their respective subarray. In the end, the edges of all vertices are extracted from their hash array.
@@
Gilbert et al. [19] tackle these problems by splitting the refinement into two parts, first unconstrained label propagation and afterwards rebalancing. The unconstrained label propagation starts by determining for each vertex $v$ the gain $F(v)$ (decrease in edge-cut) of moving it into an adjacent block. All potential vertex moves are then passed through the first filter $F(v)\geq 0\lor-F(v)<\lfloor c\cdot\operatorname{conn}(v,\Pi(v))\rfloor$. Here, $\operatorname{conn}(v,b)$ is the sum of edge weights from vertex $v$ to block $b$. To pass the first filter a move needs to be non-negative (first condition) or its increase in edge-cut must be comparatively small (second condition). The constant $c\in[0,1]$ controls the number of negative gain moves, with larger values allowing more moves. These negative moves are allowed to escape local minima. All moves that pass the first filter are collected in a list $\mathcal{X}$, which is implicitly sorted in decreasing order by gain. For each vertex $v\in\mathcal{X}$, the algorithm then recomputes the gain $\mathbb{F}(v)$ under an approximate future partition state: if a neighbor $u$ appears earlier in the ordering, it is assumed to have already moved to its new block. In a second filter, only the vertices $v\in\mathcal{X}$ with a non-negative gain $\mathbb{F}(v)$ are moved to their respective destination. To prevent oscillation each vertex that was moved is locked in the next iteration of unconstrained label propagation.
@@
Rebalancing. Since the unconstrained label propagation phase ignores the balance constraint, an additional rebalancing phase is required. Jet implements two similar strategies, weak and strong rebalancing. Weak rebalancing determines for each vertex in an overloaded block $A$ the loss of moving it to a valid destination block. Valid destination blocks are blocks $B$ with $c(B)\leq\sigma\leq L_{\max}$. The upper bound $\sigma$ is used instead of $L_{\max}$ to create a small buffer zone, so a block is not overloaded after only one move. For each overloaded block $A$ its vertices are approximately sorted increasingly by loss in a list $L_{A}$. Then a prefix $L_{A}^{*}\subseteq L_{A}$ is determined such that moving all vertices in $L_{A}^{*}$ will balance block $A$. This minimizes the loss to balance a specific block, however, the destination blocks might become overloaded, so another iteration might be necessary. Strong rebalancing takes extra care by preventing destination block from becoming overloaded. If this would happen, the vertices are moved to another underloaded block. However, these blocks must not necessarily be connected to the vertex. This results in a much greater loss but the partition will most likely be balanced.
@@
In total, Jet combines its unconstrained label propagation and rebalancing in multiple iterations and keeps the best solution encountered. The algorithm offers comparable solution quality to the CPU-based graph partitioner Metis [28], while being up to 20 times faster. Other GPU graph partitioners are Goodarzi et al. [20], Sphynx [1] and G-kway [32].
3.2 Process Mapping